OpenStudy (anonymous):

Write the direct variation equation for the following: y varies directly with x and y = 60 when x = 10. A. y = 1/6 x B. y = -1/6 x C. y = 6x D. y = -6x

OpenStudy (anonymous):

y= 60 x=10 y=6(10)=6x

OpenStudy (anonymous):

When you hear y "varies directly with" x, you should make a ratio: y to x, or y:x, or y/x. I prefer to write them as fractions for this question. 60 to 10 or 60/10 Reduce: 60/10 = 6/1 = 6 The answer is 6. -Your uncle tommy

OpenStudy (anonymous):

thanks that what I was thinking

OpenStudy (anonymous):

C